Hi. Here is my algebra question
Find three consecutive odd integers such that three times the middle integer is one more than the sum of the first and third.
Please help. I am clueless.
Thanks for your help.
Kim Goray

Answer by nerdybill(7384)

Find three consecutive odd integers such that three times the middle integer is one more than the sum of the first and third.
.
All you have to realize is that "odd integers" are a distance of 2 apart.
.
Let x = first of three odd integers
then
x+2 = middle odd integer
x+4 = third odd integer
.
then from: "three times the middle integer is one more than the sum of the first and third" we get
3(x+2) = (x + x+4) + 1
3x+6 = (2x+4) + 1
3x+6 = 2x+5
x+6 = 5
x = -1 (first odd integer)
.
middle odd integer:
x+2 = -1+2 = 1
.
third odd integer:
x+4 = -1+4 = 3
.
Answer: -1, 1, 3